Hello I have a couple questions

A distribution of raw scores with respect to x has a mean (x̅ x) of 55 and a standard deviation (sx) of 4. Convert a raw score (x) of 50 into a z score from this distribution. zx =

a. 1 b. -1.25 c. 1.25 d. -2.50 e. 2.50

2. With respect to any distribution of standard scores (a non-normal or a normal distribution of z scores), the mean of the distribution is equal to 0 and the standard deviation of the distribution is equal to 1.
a. true

b. false
Answer #1

1.

the PDF of normal distribution is = 1/σ * √2π * e ^ -(x-u)^2/ 2σ^2
standard normal distribution is a normal distribution with a,
mean of 0,
standard deviation of 1
equation of the normal curve is ( Z )= x - u / sd ~ N(0,1)
mean ( u ) = 55
standard Deviation ( sd )= 4
sample mean x =50
Z score = 50-55/4
Z score = -1.25
Answer:
Z score =-1.25
option:B

2.
True
With respect to any distribution of standard scores (a non-normal or a normal distribution of z scores)
option:A
the PDF of normal distribution is = 1/σ * √2π * e ^ -(x-u)^2/ 2σ^2
standard normal distribution is a normal distribution with a,
mean of 0,
standard deviation of 1
equation of the normal curve is ( Z )= x - u / sd ~ N(0,1)
mean ( u ) = 0
standard Deviation ( sd )= 1